can you pour the water from a full 2 liter bottle into a 2 quart pitcher without spilling anything?
Answer by TimothyLamb(4379)

1 liter = 1.05669 quarts
q/L = 1.05669/1
q/2 = 1.05669/1
q = 2*1.05669/1
q = 2.11338
---
answer:
you cannot, you'll spill 0.11338 quarts
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---